OpenStudy (anonymous):

Which of the following are measurements of the sides of a right triangle? A. 3, 4, 5 B. 28, 26, 12 C. 17, 14, 6 D. all of the above

OpenStudy (anonymous):

To find out, put the numbers through the Pythagerean Theorm. If it's ends up true, it's a right triangle: a^2+b^2=c^2 3^2+4^2=5^2 9+16=25 25=25 This one is correct.